# Towards Understanding the Mechanisms of Classifier-Free Guidance

### v1

@@ B.2 CFG-Guided ODE
where $-f(\sigma)c_{k}(\sigma)d\sigma$ corresponds to the combination of the standard conditional score (36) and the CPC guidance term (37). Integrating over both sides of (52), we get:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
We validate Theorem 1 on a 2D toy model with $\bm{U}_{c}=\left[\begin{smallmatrix}\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{2}}\\ \frac{1}{\sqrt{2}}&\frac{-1}{\sqrt{2}}\end{smallmatrix}\right]$, $\bm{\Lambda}_{c}=\left[\begin{smallmatrix}10&0\\ 0&3\end{smallmatrix}\right]$ and $\bm{\Lambda}_{uc}=\left[\begin{smallmatrix}3&0\\ 0&10\end{smallmatrix}\right]$. Figure 10 shows the effects of CFG under different class mean $\bm{\mu}_{c}$ (with $\gamma=1$ and $\bm{\mu}_{uc}=\bm{0}$). As predicted, CFG enhances variation along the positive CPC $\left[\begin{smallmatrix}\frac{1}{\sqrt{2}}&\frac{1}{\sqrt{2}}\end{smallmatrix}\right]^{T}$, suppresses variation along the negative CPC $\left[\begin{smallmatrix}\frac{1}{\sqrt{2}}&-\frac{1}{\sqrt{2}}\end{smallmatrix}\right]^{T}$, and shifts samples roughly toward $\bm{\mu}_{c}-\bm{\mu}_{uc}$ at a rate proportional to $\gamma||\bm{\mu}_{c}-\bm{\mu}_{uc}||_{2}$.
@@ Appendix D Naive Conditional Generation Lacks Class-Specificity
In section 3.1 we argue that naive conditional generation lacks class-specificity and in the linear model setting, such issue can be partially attributed to the non-distinctiveness of the class covariance matrices. In this section, we provide comprehensive experiments to support our claim both qualitatively and quantitatively.
@@ D.1 Qualitative Results
We generate samples using naive conditional sampling (4) and CFG sampling (6) for all 10 classes of CIFAR-10, as well as for 10 selected ImageNet classes: including (i) class 0: tench , (ii) class 31: tree frog, (iii) class 64: green mamba, (iv) class 207: golden retriever, (v) class 430: basketball, (vi) class 483: castle, (vii) class 504: coffee mug, (viii) class 817: sports car, (ix) class 933: cheese burger and (x) class 947: mushroom. CFG is applied to the entire noise interval $\sigma(t)\in[0.002,80]$, with guidance strength $\gamma=4$. The results for CIFAR-10 and ImageNet are shown in Figure 11 and Figure 12 respectively.
@@
Figure 12: Effects of CFG on ImageNet. (a) and (b) demonstrate the naive conditional samples and the CFG-guided samples of deep diffusion model and linear diffusion model respectively. Each grid corresponds to the same initial noise. Here we only display 6 classes since the other 4 classes are presented in fig. 2.
@@ D.2 Quantitative Results
  • Classifier-free guidance (CFG). $X$ and $Y$ contain images generated from the same EDM model using CFG sampling (6). We generate approximately the same number of images as the corresponding training images.
The results are presented in Figure 13, which shows that for most pairs of classes, when $X$ and $Y$ are built with images generated with naive conditional sampling, the FID (colored in orange) is consistently lower than when they are built with training data (colored in blue). Because lower FID indicates higher similarity, this results confirms that images produced by naive conditional sampling are less distinguishable across classes than the real data. In contrast, CFG greatly improves the FID score, implying an increased inter-class separation.
The samples used for calculating FID in Figure 13 are generated using 20 steps of Euler method (first-order sampler). Increasing the number of steps or switching to higher-order sampler only marginally narrows the gap. Table 1 shows the inter-class FID averaged over 10 selected classes as described in section D.1 with different sampling steps and sampler. Note that even when using 100 steps and second-order Heun samplers, the average inter-class FID is still considerably smaller compared to the training data (ground truth). Figure 14 qualitatively visualizes the samples generated from the same initial noise but different class labels. Despite conditioned on different labels, the generated images share high structural similarity. For certain classes, such as tree frog, green mamba and golden retriever, the class features are even hard to discern. In contrast, CFG greatly reduces the structural similarity, yielding images with clear, class-specific features.
@@
Table 1: Average inter-class FID for training data and various sampling settings (10-class average).
@@ D.3 Covariance Matrices of Different Classes Lack Class-Specificity
The lack of class-specificity is especially pronounced in linear diffusion models. As shown in Figure 14(a) and Figure 15, although the linear diffusion models are separately parameterized with the class-specific means and covariances for each class, the resulting samples share high similarity. Since the generated samples of the linear models are governed by the data covariances, the observed inter-class similarity implies that the covariance structures of different classes are not distinct enough.
Next, we quantitatively demonstrate that the class-specific covariance matrices are insufficiently distinct. To do this, we take $\bm{U}_{uc}$, the principal components (PCs) of the unconditional dataset (i.e., the singular vectors of the unconditional covariance), as a baseline. We then compare $\bm{U}_{uc}$ to $\bm{U}_{c}$, the PCs of each conditional dataset. As shown in Figure 16 and Figure 17, the correlation matrices $\bm{U}_{c}^{T}\bm{U}_{uc}$ for 10 classes (5 from CIFAR-10 and 5 from ImageNet) reveal that the leading PCs of each class share high similarity with those of the unconditional data. Thus, the PCs do not necessarily capture the distinctive features of individual classes though they represent the dominant variations of the dataset. Instead, these PCs often reflect global intensity or foreground-background variations.
@@ E.1 Mean-Shift Guidance
Equation 38 is the mean-shift guidance term that shifts $\bm{x}_{t}$ towards $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})$, a direction independent of the specific sample $\bm{x}_{t}$. At sufficiently large $\sigma(t)$, $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})\approx\bm{\mu}_{c}-\bm{\mu}_{uc}$, indicating the mean-shift term approximately shifts $\bm{x}_{t}$ towards the direction of the difference between class mean and unconditional mean $\bm{\mu}_{c}-\bm{\mu}_{uc}$. As $\sigma(t)$ decreases, the components of $\bm{\mu}_{c}-\bm{\mu}_{uc}$ within the subspace spanned by the unconditional PCs ($\bm{U}_{uc}$) progressively shrink to 0. Figure 18 demonstrates $\bm{\mu}_{c}-\bm{\mu}_{uc}$ and the evolution of the mean-shift guidance term $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})$ across different noise levels. Notice that for a wide range of noise levels $\sigma(t)$, $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})$ remains close to $\bm{\mu}_{c}-\bm{\mu}_{uc}$, before it becomes uninformative at small $\sigma(t)$. Hence, as stated in the main text, the mean-shift guidance term has the effect of approximately shifting $\bm{x}_{t}$ in the direction $\bm{\mu}_{c}-\bm{\mu}_{uc}$.
@@ E.2 CPC guidance
Equation 37 is the CPC guidance term. Let $\bm{V}_{\sigma(t)}\hat{\bm{\Lambda}}_{\sigma(t)}\bm{V}_{\sigma(t)}^{T}$ be the eigendecomposition of $\tilde{\bm{\Sigma}}_{c,t}-\tilde{\bm{\Sigma}}_{uc,t}$, whose eigen spectrum is demonstrated in Figure 19, the CPC guidance term can be further decomposed into the positive CPC and negative CPC guidance:
@@
where $\bm{V}_{\sigma(t),+}$ and $\bm{V}_{\sigma(t),-}$ contain eigenvectors corresponding to positive and negative eigenvalues $\hat{\bm{\Lambda}}_{\sigma(t),+}$ and $\hat{\bm{\Lambda}}_{\sigma(t),-}$ respectively. As discussed in section 2.3, $\tilde{\bm{\Sigma}}_{c,t}$ and $\tilde{\bm{\Sigma}}_{uc,t}$ are (up to a factor $\sigma(t)^{2}$) the conditional and unconditional posterior covariances of $p_{\text{data}}(\bm{x}|\bm{c})=\mathcal{N}(\bm{\mu}_{c},\bm{\Sigma}_{c})$ and $p_{\text{data}}(\bm{x})=\mathcal{N}(\bm{\mu}_{uc},\bm{\Sigma}_{uc})$. Hence, $\bm{V}_{\sigma(t)}$ are the CPCs which contrast between $X\sim p_{data}(\bm{x}|\bm{x}_{t},\bm{c})$ and $Y\sim p_{\text{data}}(\bm{x}|\bm{x}_{t})$. Specifically, $\bm{V}_{\sigma(t),+}$ captures directions of higher conditional variance (class-specific features), while $\bm{V}_{\sigma(t),-}$ captures directions of higher unconditional variance (features more relevant to the unconditonal data). Figure 20 illustrates the evolution of positive CPCs ($\bm{V}_{\sigma(t),+}$) and PCs ($\bm{U}_{c}$) across different noise levels. It is evident that the positive CPCs better capture the class-specific patterns compared to PCs. Here we choose not to display negative CPCs since they correspond to generic features that explain more variances for the unconditional dataset, which are less visually interpretable. Nevertheless, as we will show next, suppressing these directions is beneficial.
@@ E.3 Distinct Effects of the CFG Components
Figures 21 and 22 qualitatively demonstrates the effects of each CFG component in linear diffusion models over 10 different ImageNet classes.
@@
As described in section B.3, the distortion effects of the CFG components can be quantitatively verified through the following experiment:
@@
The resulting univariate distributions quantify the amount of energy the samples have along these directions. The above experiment are performed on both linear and nonlinear (EDM) diffusion models. The samples are generated using 20 steps and the guidance strength $\gamma$ is set to 2. We focus on the first class of ImageNet (tench) and present the results on the first 5 positive CPCs and negative CPCs. As shown in Figures 23 and 24, compared to the samples with no CFG, the distributions of the CFG-guided samples have higher density on the positive CPC directions but lower density on the negative CPC directions, implying the former is enhanced while the latter is suppressed. The univariate distribution of the projection onto the mean-shift direction is presented in Figure 4(c) (bottom row), from which it is clear that the density is shifted in the direction of $\bm{\mu}_{c}-\bm{\mu}_{uc}$.
@@ Appendix F CFG in Nonlinear Deep Diffusion Models
In this section we provide additional experimental results for section 4, where we investigate how well the insights derived from linear diffusion models extend to real-world, nonlinear deep diffusion models. In this work, we study the state-of-the-art EDM models [4].
@@ F.1 Linear to Nonlinear Transition in Diffusion Models
Let $\bm{U}_{t}(\bm{x}_{t})$ be the left singular vectors of the network Jacobians $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ along the sampling trajectories, and let $\bm{U}_{c}$ be the left singular vectors of $\nabla\mathcal{D}_{\mathrm{L}}(\bm{x}_{t};\sigma(t),\bm{c})$. Since $\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))=\bm{\mu}_{c}+\bm{U}_{c}\tilde{\bm{\Lambda}}_{c,\sigma(t)}\bm{U}_{c}^{T}(\bm{x}-\bm{\mu}_{c})$, if $\mathcal{D}_{\bm{\theta}}\approx\mathcal{D}_{\mathrm{L}}$, then $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})\approx\nabla\mathcal{D}_{\mathrm{L}}(\bm{x}_{t};\sigma(t),\bm{c})\approx\bm{U}_{c}\tilde{\bm{\Lambda}}_{c,\sigma(t)}\bm{U}_{c}^{T}$, implying $\bm{U}_{t}(\bm{x}_{t})\approx\bm{U}_{c}$, independent of $\bm{x}_{t}$. As illustrated in Figure 25, for large $\sigma(t)$, the leading singular vectors of $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ indeed align with $\bm{U}_{c}$. Note that since $\tilde{\bm{\Lambda}}_{c,\sigma(t)}=\text{diag}(\frac{\lambda_{c,1}}{\lambda_{c,1}+\sigma^{2}(t)},...,\frac{\lambda_{c,d}}{\lambda_{c,d}+\sigma^{2}(t)})$, $\nabla\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))$ is highly low-rank at large $\sigma(t)$. Thus, our primary interest is in the leading singular vectors, and the non-leading singular vectors are ambiguous. In contrast, for small $\sigma(t)$, the alignment no longer holds and $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t))$ starts to adapt to individual samples, reflecting the model’s nonlinear behavior. Figures 5, 26 and 27 qualitatively demonstrates this linear to nonlinear transition.

### v3

@@ B.2 CFG-Guided ODE
where $f(\sigma)c_{k}(\sigma)d\sigma$ corresponds to the combination of the standard conditional score (36) and the CPC guidance term (37). Integrating over both sides of (52), we get:
@@
$\int-f(\sigma)d\sigma$ $=\int\frac{\sigma}{\lambda_{c,k}+\sigma^{2}}d\sigma-\gamma\int\frac{\sigma(\lambda_{c,k}-\lambda_{uc,k})}{(\lambda_{c,k}+\sigma^{2})(\lambda_{uc,k}+\sigma^{2})}d\sigma$ (54)
@@
$c_{k}(\sigma)=\hat{c}_{k}(\sigma)+\frac{1}{I(\sigma)}\int I(\sigma^{\prime})g(\sigma^{\prime})d\sigma^{\prime},$ (61)
where $I(\sigma)=e^{\int f(\sigma^{\prime})d\sigma^{\prime}}$ is the integrating factor. Since:
@@
$c_{k}(\sigma)$ $=\hat{c}_{k}(\sigma)+\gamma(\lambda_{c,k}+\sigma^{2})^{\frac{1}{2}}(\frac{\lambda_{c,k}+\sigma^{2}}{\lambda_{uc,k}+\sigma^{2}})^{\frac{\gamma}{2}}\int_{\sigma}^{\sigma_{T}}\frac{(\lambda_{uc,k}+\tilde{\sigma}^{2})^{\frac{\gamma}{2}-1}}{(\lambda_{c,k}+\tilde{\sigma}^{2})^{\frac{\gamma+1}{2}}}\bm{u}_{c,k}^{T}(\bm{\mu}_{c}-\bm{\mu}_{uc})\tilde{\sigma}d\tilde{\sigma}$ (63)
@@
where $b_{\sigma,k}=(\lambda_{c,k}+\sigma^{2})^{\frac{1}{2}}(\frac{\lambda_{c,k}+\sigma^{2}}{\lambda_{uc,k}+\sigma^{2}})^{\frac{\gamma}{2}}\int_{\sigma}^{\sigma_{T}}\frac{(\lambda_{uc,k}+\tilde{\sigma}^{2})^{\frac{\gamma}{2}-1}}{(\lambda_{c,k}+\tilde{\sigma}^{2})^{\frac{\gamma+1}{2}}}\tilde{\sigma}d\tilde{\sigma}$. Therefore we have:
@@
[…]
@@ Appendix D Naive Conditional Generation Lacks Class-Specificity
In section˜3.1 we argue that naive conditional generation lacks class-specificity and in the linear model setting, such issue can be partially attributed to the non-distinctiveness of the class covariance matrices. In this section, we provide comprehensive experiments to support our claim both qualitatively and quantitatively.
@@ D.1 Qualitative Results
We generate samples using naive conditional sampling (4) and CFG sampling (6) for all 10 classes of CIFAR-10, as well as for 10 selected ImageNet classes: including (i) class 0: tench , (ii) class 31: tree frog, (iii) class 64: green mamba, (iv) class 207: golden retriever, (v) class 430: basketball, (vi) class 483: castle, (vii) class 504: coffee mug, (viii) class 817: sports car, (ix) class 933: cheese burger and (x) class 947: mushroom. CFG is applied to the entire noise interval $\sigma(t)\in[0.002,80]$, with guidance strength $\gamma=4$. The results for CIFAR-10 and ImageNet are shown in Figure˜11 and Figure˜12 respectively.
@@
Figure 12: Effects of CFG on ImageNet. (a) and (b) demonstrate the naive conditional samples and the CFG-guided samples of deep diffusion model and linear diffusion model respectively. Each grid corresponds to the same initial noise. Here we only display 6 classes since the other 4 classes are presented in fig.˜2.
@@ D.2 Quantitative Results
  • Classifier‑free guidance (CFG). $X$ and $Y$ contain images generated from the same EDM model using CFG sampling (6). We generate approximately the same number of images as the corresponding training images.
The results are presented in Figure˜13, which shows that for most pairs of classes, when $X$ and $Y$ are built with images generated with naive conditional sampling, the FID (colored in orange) is consistently lower than when they are built with training data (colored in blue). Because lower FID indicates higher similarity, this results confirms that images produced by naive conditional sampling are less distinguishable across classes than the real data. In contrast, CFG greatly improves the FID score, implying an increased inter-class separation.
The samples used for calculating FID in Figure˜13 are generated using 20 steps of Euler method (first-order sampler). Increasing the number of steps or switching to higher-order sampler only marginally narrows the gap. Table˜1 shows the inter-class FID averaged over 10 selected classes as described in section˜D.1 with different sampling steps and sampler. Note that even when using 100 steps and second-order Heun samplers, the average inter-class FID is still considerably smaller compared to the training data (ground truth). Figure˜14 qualitatively visualizes the samples generated from the same initial noise but different class labels. Despite conditioned on different labels, the generated images share high structural similarity. For certain classes, such as tree frog, green mamba and golden retriever, the class features are even hard to discern. In contrast, CFG greatly reduces the structural similarity, yielding images with clear, class-specific features.
@@
Table 1: Average inter‑class FID for training data and various sampling settings (10‑class average).
@@ D.3 Covariance Matrices of Different Classes Lack Class-Specificity
The lack of class-specificity is especially pronounced in linear diffusion models. As shown in Figure˜14(a) and Figure˜15, although the linear diffusion models are separately parameterized with the class-specific means and covariances for each class, the resulting samples share high similarity. Since the generated samples of the linear models are governed by the data covariances, the observed inter-class similarity implies that the covariance structures of different classes are not distinct enough.
Next, we quantitatively demonstrate that the class-specific covariance matrices are insufficiently distinct. To do this, we take $\bm{U}_{uc}$, the principal components (PCs) of the unconditional dataset (i.e., the singular vectors of the unconditional covariance), as a baseline. We then compare $\bm{U}_{uc}$ to $\bm{U}_{c}$, the PCs of each conditional dataset. As shown in Figure˜16 and Figure˜17, the correlation matrices $\bm{U}_{c}^{T}\bm{U}_{uc}$ for 10 classes (5 from CIFAR-10 and 5 from ImageNet) reveal that the leading PCs of each class share high similarity with those of the unconditional data. Thus, the PCs do not necessarily capture the distinctive features of individual classes though they represent the dominant variations of the dataset. Instead, these PCs often reflect global intensity or foreground-background variations.
@@ E.1 Mean-Shift Guidance
Equation˜38 is the mean-shift guidance term that shifts $\bm{x}_{t}$ towards $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})$, a direction independent of the specific sample $\bm{x}_{t}$. At sufficiently large $\sigma(t)$, $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})\approx\bm{\mu}_{c}-\bm{\mu}_{uc}$, indicating the mean-shift term approximately shifts $\bm{x}_{t}$ towards the direction of the difference between class mean and unconditional mean $\bm{\mu}_{c}-\bm{\mu}_{uc}$. As $\sigma(t)$ decreases, the components of $\bm{\mu}_{c}-\bm{\mu}_{uc}$ within the subspace spanned by the unconditional PCs ($\bm{U}_{uc}$) progressively shrink to 0. Figure˜18 demonstrates $\bm{\mu}_{c}-\bm{\mu}_{uc}$ and the evolution of the mean-shift guidance term $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})$ across different noise levels. Notice that for a wide range of noise levels $\sigma(t)$, $(\bm{I}-\tilde{\bm{\Sigma}}_{uc,t})(\bm{\mu}_{c}-\bm{\mu}_{uc})$ remains close to $\bm{\mu}_{c}-\bm{\mu}_{uc}$, before it becomes uninformative at small $\sigma(t)$. Hence, as stated in the main text, the mean-shift guidance term has the effect of approximately shifting $\bm{x}_{t}$ in the direction $\bm{\mu}_{c}-\bm{\mu}_{uc}$.
@@ E.2 CPC guidance
Equation˜37 is the CPC guidance term. Let $\bm{V}_{\sigma(t)}\hat{\bm{\Lambda}}_{\sigma(t)}\bm{V}_{\sigma(t)}^{T}$ be the eigendecomposition of $\tilde{\bm{\Sigma}}_{c,t}-\tilde{\bm{\Sigma}}_{uc,t}$, whose eigen spectrum is demonstrated in Figure˜19, the CPC guidance term can be further decomposed into the positive CPC and negative CPC guidance:
@@
where $\bm{V}_{\sigma(t),+}$ and $\bm{V}_{\sigma(t),-}$ contain eigenvectors corresponding to positive and negative eigenvalues $\hat{\bm{\Lambda}}_{\sigma(t),+}$ and $\hat{\bm{\Lambda}}_{\sigma(t),-}$ respectively. As discussed in section˜2.3, $\tilde{\bm{\Sigma}}_{c,t}$ and $\tilde{\bm{\Sigma}}_{uc,t}$ are (up to a factor $\sigma(t)^{2}$) the conditional and unconditional posterior covariances of $p_{\text{data}}(\bm{x}|\bm{c})=\mathcal{N}(\bm{\mu}_{c},\bm{\Sigma}_{c})$ and $p_{\text{data}}(\bm{x})=\mathcal{N}(\bm{\mu}_{uc},\bm{\Sigma}_{uc})$. Hence, $\bm{V}_{\sigma(t)}$ are the CPCs which contrast between $X\sim p_{data}(\bm{x}|\bm{x}_{t},\bm{c})$ and $Y\sim p_{\text{data}}(\bm{x}|\bm{x}_{t})$. Specifically, $\bm{V}_{\sigma(t),+}$ captures directions of higher conditional variance (class-specific features), while $\bm{V}_{\sigma(t),-}$ captures directions of higher unconditional variance (features more relevant to the unconditonal data). Figure˜20 illustrates the evolution of positive CPCs ($\bm{V}_{\sigma(t),+}$) and PCs ($\bm{U}_{c}$) across different noise levels. It is evident that the positive CPCs better capture the class-specific patterns compared to PCs. Here we choose not to display negative CPCs since they correspond to generic features that explain more variances for the unconditional dataset, which are less visually interpretable. Nevertheless, as we will show next, suppressing these directions is beneficial.
@@ E.3 Distinct Effects of the CFG Components
Figures˜21 and 22 qualitatively demonstrates the effects of each CFG component in linear diffusion models over 10 different ImageNet classes.
@@
The distortion effects of the CFG components can be quantitatively verified through the following experiment:
@@
The resulting univariate distributions quantify the amount of energy the samples have along these directions. The above experiment are performed on both linear and nonlinear (EDM) diffusion models. The samples are generated using 20 steps and the guidance strength $\gamma$ is set to 2. We focus on the first class of ImageNet (tench) and present the results on the first 5 positive CPCs and negative CPCs. As shown in Figures˜23 and 24, compared to the samples with no CFG, the distributions of the CFG-guided samples have higher density on the positive CPC directions but lower density on the negative CPC directions, implying the former is enhanced while the latter is suppressed. The univariate distribution of the projection onto the mean-shift direction is presented in Figure˜4(c) (bottom row), from which it is clear that the density is shifted in the direction of $\bm{\mu}_{c}-\bm{\mu}_{uc}$.
@@ Appendix F CFG in Nonlinear Deep Diffusion Models
In this section we provide additional experimental results for section˜4, where we investigate how well the insights derived from linear diffusion models extend to real-world, nonlinear deep diffusion models. In this work, we study the state-of-the-art EDM models [4].
@@ F.1 Linear to Nonlinear Transition in Diffusion Models
Let $\bm{U}_{t}(\bm{x}_{t})$ be the left singular vectors of the network Jacobians $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ along the sampling trajectories, and let $\bm{U}_{c}$ be the left singular vectors of $\nabla\mathcal{D}_{\mathrm{L}}(\bm{x}_{t};\sigma(t),\bm{c})$. Since $\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))=\bm{\mu}_{c}+\bm{U}_{c}\tilde{\bm{\Lambda}}_{c,\sigma(t)}\bm{U}_{c}^{T}(\bm{x}-\bm{\mu}_{c})$, if $\mathcal{D}_{\bm{\theta}}\approx\mathcal{D}_{\mathrm{L}}$, then $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})\approx\nabla\mathcal{D}_{\mathrm{L}}(\bm{x}_{t};\sigma(t),\bm{c})\approx\bm{U}_{c}\tilde{\bm{\Lambda}}_{c,\sigma(t)}\bm{U}_{c}^{T}$, implying $\bm{U}_{t}(\bm{x}_{t})\approx\bm{U}_{c}$, independent of $\bm{x}_{t}$. As illustrated in Figure˜25, for large $\sigma(t)$, the leading singular vectors of $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$ indeed align with $\bm{U}_{c}$. Note that since $\tilde{\bm{\Lambda}}_{c,\sigma(t)}=\text{diag}(\frac{\lambda_{c,1}}{\lambda_{c,1}+\sigma^{2}(t)},...,\frac{\lambda_{c,d}}{\lambda_{c,d}+\sigma^{2}(t)})$, $\nabla\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))$ is highly low-rank at large $\sigma(t)$. Thus, our primary interest is in the leading singular vectors, and the non-leading singular vectors are ambiguous. In contrast, for small $\sigma(t)$, the alignment no longer holds and $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t))$ starts to adapt to individual samples, reflecting the model’s nonlinear behavior. Figures˜5, 26 and 27 qualitatively demonstrates this linear to nonlinear transition.